OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <10, 0>, v = <0, -9> will give medals

OpenStudy (freckles):

Have you computed the dot product?

OpenStudy (freckles):

If the dot product is 0 then u and v are orthogonal if you don't get 0 you have more work to determine if they are parallel or not... they are parallel if the angle between them is either 0 degrees or 180 degrees

OpenStudy (anonymous):

how do you get the dot product?

OpenStudy (freckles):

\[<a,b> \cdot <c,d>=a \times c+b \times d\]

OpenStudy (anonymous):

so its 0

OpenStudy (freckles):

yes 10(0)+0(-9) is 0+0 which is 0

OpenStudy (freckles):

any questions on this question before I go

OpenStudy (anonymous):

ohh okay thank you soo much no questions

OpenStudy (freckles):

k cool